40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 16:46:31 2019, gautam, Update, IOO, IMC not locking PCdrive_RMS.png
    Reply  Tue Jun 4 10:56:31 2019, gautam, Update, IOO, IMC diagnostics IMCdiag.pdf
       Reply  Thu Jun 6 22:11:53 2019, Koji, Update, IOO, IMC diagnostics 
          Reply  Thu Jun 13 18:01:18 2019, aaron, Update, IOO, IMC diagnostics 
             Reply  Fri Jun 14 13:10:00 2019, aaron, Update, IOO, IMC diagnostics 
                Reply  Mon Jun 17 12:37:16 2019, aaron, Update, IOO, IMC diagnostics 
                   Reply  Tue Jun 18 20:35:07 2019, aaron, Update, IOO, IMC diagnostics 
                      Reply  Wed Jun 19 19:12:51 2019, aaron, Update, IOO, IMC diagnostics PMC_Error_Spectrum.pdfPDH_signal.pdfPDH_signal_full.pdf
                         Reply  Mon Jun 24 15:49:05 2019, aaron, Update, IOO, IMC diagnostics sweep_config_updated.png
                            Reply  Tue Jul 2 19:36:18 2019, aaron, Update, IOO, IMC diagnostics PDH_error.pdfPMC_Error_Spectrum.pdf
          Reply  Fri Jun 21 19:36:26 2019, Koji, Update, IOO, IMC diagnostics IMC_error.pdf
             Reply  Sun Jun 23 08:09:53 2019, gautam, Update, IOO, NPRO diagnostics consolidatedOscopeScreenCaps.pdfRINcomp.pdfPLL_OLTF.pdfPLLnoise.pdf
                Reply  Sun Jun 23 09:36:32 2019, gautam, Update, IOO, IMC is locking normally again PMCdiag.pdf
                   Reply  Sun Jun 23 14:43:14 2019, Koji, Update, IOO, IMC is locking normally again 
                   Reply  Mon Jun 24 08:12:10 2019, gautam, Update, IOO, IMC is locking normally again IMCdutycycle.png
                      Reply  Mon Jun 24 11:48:35 2019, gautam, Update, IOO, IMC in-loop error spectra and OLTF IMCfreqNoise.pdfIMC_OLTF.pdfIMC_CMboard.jpg
Message ID: 14675     Entry time: Fri Jun 14 13:10:00 2019     In reply to: 14670     Reply to this: 14677
Author: aaron 
Type: Update 
Category: IOO 
Subject: IMC diagnostics 

The circuit diagram for the PMC servo card is D980352. From this diagram, I see that I can send an excitation from the network analyzer to FP2TEST (9.09 kOhm input impedance) where it is added to the PMC error signal before going to the loop filters.

I hook up the following

  • Agilent 4395A output to SR560 (300 Hz HP, gain of 1)
  • SR560 to FP2TEST and to Agilent's channel R
  • PMC error signal IF (mixer box mounted to rack, I noticed the IF BNC->SMA were a bit loose/stressed by a hanging LP RF filter) to SR560 (also 300 Hz HP, gain of 2)
  • SR560 to Agilent channel A

I 'Enable' Test 2 on the PSL screen, so FP2TEST gets added to the error signal.

PDH signal

  • TDS 3034B with four channels
    • 1. PMC servo box external drive (split off from the function generator)
    • 2. PMC servo box output monitor (mirrors the drive, shows when drive is saturating)
    • 3. IF signal (split off after the mixer)
    • 4. PMC Trans (long BNC from the PSL table)
  • SRS DS345 function generator (into the PMC servo box' external drive)
    • 100 Hz signal (there's a 10 Hz pole on the PZT drive, so any faster than this and I can't see both sidebands without the HV output mon clipping)
    • 3.19 Vpp amplitude (smallest amplitude at 100 Hz such that both sidebands are well resolved)
    • 1.52 V offset (center the carrier's PDH error signal at pi/2 out of phase with the drive)

I was able to see the carrier and both sidebands.

I tried to grab this data from the scope via ethernet, but was unsuccessful (timeout errors, I'm using the scripts from scripts/tektronix/tek-dump, and the GPIB box that Kruthi had been using for the GigE cam; I also tried plugging in directly scope->ethernet. Never got anything but timeout errors, so maybe I'm not specifying the port correctly. Anyway the trace is frozen on the scope for later use, or I can easily repeat this now that I know how).


Next, I locked the PMC (Test1 is off, tune DC output adjust until I get some transmission, turn on the loop at Test1, increase the gain to before the loop goes unstable). I'm sending the following channels to SR560 (gain = 2, no filtering, high dynamic reserve, 50 Ohm outputs), and reading spectra from the Agilent 4395A:

  • A-- PDH IF

The HV mon was always saturating the preamp, so I disconnected it; I added a 50 Hz (6db) high pass to the Trans PD signal, since it has a DC component.

I got to take a look at the traces on the spectrum analyzer front panel, but too tired to do the GPIB for now. There are peaks, things look reasonable.

ELOG V3.1.3-